OpenStudy (anonymous):

Solve 4^x=20 Round to the nearest ten thousandth

OpenStudy (lgbasallote):

do you know how to convert it to logarithmic form?

OpenStudy (anonymous):

No.. my teacher sucks. :)

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

take log number on log the base and it will be x=log20/log4

OpenStudy (anonymous):

\[4^x=20\] \[x \log 4 = \log 20\] \[x=\frac{\log20}{\log 4}\] Basically, log both sides. Usually the log rule, you can take the exponent x and place it as a coefficient. Then divide both sides to solve for x.

OpenStudy (anonymous):

Okay

OpenStudy (lgbasallote):

lol when i said change to log form i meant \[\LARGE a^x = b \Longleftrightarrow \log_a b = x\]

OpenStudy (anonymous):

Ohh I think I learned that. I think. lol

OpenStudy (lgbasallote):

well you end up with \[\large \frac{\log 20}{\log 4}\] anyway..have you learned that?

OpenStudy (anonymous):

yea.. I think so. But we learn stuff so quick i never get a chance to fully understand.

OpenStudy (anonymous):

I got the answer now :)